OpenStudy (anonymous):

Quadrilateral ABCD is a parallelogram. The coordinates are A (2,4), B (4,7), C (5,1). What is the coordinate of D? I've tried graphing but I wish to do this algebraically. How should I start?

OpenStudy (anonymous):

hi .. Here under simple sketch for your problem .. I'm sure its so easy when you draw it .  while its parallelogram ( all time mistake by right this word i keep copy paste :P ) then shift displacement between A B , the same between C D first we want to find displacement B-A = ( 2, 3 ) D = C + (B-A ) = ( 7 , 4 ) That is all !